OpenStudy (anonymous):

Which of the following represents 6x^3/4 in radical form?

OpenStudy (anonymous):

when you have fractional exponents, you can change them to radicals by moving the denominator of the fraction to be the base of the radical example x^(1/2) = \[\sqrt[2]{x^1}\] which is also equal to \[\sqrt{x}\]

OpenStudy (anonymous):

awwww you're so sweet for typing all of that up :)

OpenStudy (anonymous):

What would i put in place of x?

OpenStudy (anonymous):

depends. in the original problem (just from the way you wrote it) i would assume that its just x, and you can follow the same procedure as before, but multiply the radical by 6...it could also be the case that the problem is (6x)^(3/4), in that case, just substitute 6x in for x

OpenStudy (anonymous):

i would think its the first way. if theres no parenthesis around the 6x in the original problem, you can multiply the 4th root of x^3 by 6

OpenStudy (anonymous):

and that would be your answer

OpenStudy (anonymous):

Thank you for explaining!!!!!!!1 :))))))))

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

So is this right? :

OpenStudy (anonymous):

4 sq root 6x^3

OpenStudy (anonymous):

\[6\sqrt[4]{x ^{3}}\] its like that because if the 6x isnt in parenthesis, the exponent only gets applied to the x

OpenStudy (anonymous):

sorry...! I'm only confusing myself here LOL :) Thank you so much, you're so kind and thoughtful!!!!!!! :)

OpenStudy (anonymous):

welcome :)